x^2-9x=0
I am told to factor
Please help i have no idea!

Found 2 solutions by checkley77, Alan3354:
Answer by checkley77(12844)

x^2-9x=0
x(x-9)=0
x=0 ans.
x-9=0
x=9 ans.

Answer by Alan3354(69443)

x^2-9x=0
I am told to factor
-----------
Factoring is the first step.
x^2-9x=0
x*(x-9) = 0
Then solve for x.
x = 0, or
x = 9